OpenStudy (anonymous):

Which of the following expressions are equivalent? Justify your reasoning. A.4√x^3 B. 1 divide x−1 C. ^10√x^5•x^4•x^2 D.x^1/3•x^ 1/3•x^1/3

OpenStudy (anonymous):

\[x ^{\frac{ 1 }{ 3 }\times3}\] would be D simplified

OpenStudy (anonymous):

A would be \[x ^{\frac{ 3 }{ 4 }}\] and C would be \[x ^{\frac{ 11 }{ 10 }}\]

OpenStudy (anonymous):

@texaschic101